OpenStudy (lgbasallote):

if P(A) = 0.75 and P(B = 0.8) and P(A' \(\cap\) B') = 0.05 are events A and B independent?

OpenStudy (calculusfunctions):

Two events A and B are independent events iff P(A and B) = P(A) x P(B).

OpenStudy (lgbasallote):

hmmm

OpenStudy (lgbasallote):

that's not dependent?

OpenStudy (calculusfunctions):

Nope!

OpenStudy (calculusfunctions):

In your question, P(not A) = 0.25 and the P(not B) = 0.2 and P(not A and not B) = 0.95. P(not A) x P(not B) = 0.25 x 0.2 = 0.05 = P(not A and not B) ∴ A and B are independent events.

OpenStudy (lgbasallote):

i thought B was independent if P(A|B) = P(A'|B)

OpenStudy (lgbasallote):

nevermind. i got it

OpenStudy (calculusfunctions):

Oh my gosh! There is something majorly wrong with my computer. I was almost finished with the explanation and it froze. I'll start over. Give me a few minutes.

OpenStudy (calculusfunctions):

OK good so I don't have to start over. You see that my explanation was right.

OpenStudy (lgbasallote):

yes except for a minor typographical error but understandable

OpenStudy (calculusfunctions):

Good! Let me know if you change your mind and want to see the explanation.